# H-STAR: LLM-driven Hybrid SQL-Text Adaptive Reasoning on Tables

### v1

@@ Abstract
Tabular reasoning involves interpreting unstructured queries against structured tables, requiring a synthesis of textual understanding and symbolic reasoning. Existing methods rely on either of the approaches and are constrained by their respective limitations. Textual reasoning excels in semantic interpretation unlike symbolic reasoning (SQL logic), but falls short in mathematical reasoning where SQL excels. In this paper, we introduce a novel algorithm H-STAR, comprising table extraction and adaptive reasoning, integrating both symbolic and semantic (text-based) approaches. To enhance evidence extraction, H-STAR employs a multi-view approach, incorporating step-by-step row and column retrieval. It also adapts reasoning strategies based on question types, utilizing symbolic reasoning for quantitative and logical tasks, and semantic reasoning for direct lookup and complex lexical queries. Our extensive experiments demonstrate that H-STAR significantly outperforms state-of-the-art methods across three tabular question-answering (QA) and fact-verification datasets, underscoring its effectiveness and efficiency. [^1] The codes and data are available at: https://github.com/nikhilsab/H-STAR
@@ 1 Introduction
Tabular data is one of the most widely used formats for storing structured information in real-world applications. Table-based reasoning presents an inherently challenging problem, requiring logical, mathematical, and textual reasoning over unstructured queries and structured tables Ye et al. (2023). Thus, understanding tabular data has become a significant area of research in natural language processing. Tabular reasoning tasks (illustrated in Figure 1), such as table-based question answering Pasupat and Liang (2015); Nan et al. (2022) and table-based fact verification Chen et al. (2019), have been extensively explored in the past.
@@
Large Language Model (LLM) based prompting methods involving symbolic approaches such as Program of Thought Chen et al. (2023) and textual methods like Chain of Thought (CoT) Wei et al. (2022); Brown et al. (2020) have been used for table reasoning. As shown in Figure 2, these methods often fall short due to the complexities imposed by the intricate mix of numerical, temporal, and textual data, coupled with complex table structures Shi et al. (2023); Liu et al. (2023); Sui et al. (2024). Recent techniques Cheng et al. (2022); Ye et al. (2023); Wang et al. (2023); Nahid and Rafiei (2024) rely on either textual or symbolic reasoning. Textual reasoning excels in natural language understanding but often misinterprets table structures and struggles with quantitative reasoning. Conversely, SQL-based approaches are strong in quantitative problem-solving but perform poorly on noisy or unstructured inputs Liu et al. (2023). Integrating both methods can leverage their complementary strengths, thereby addressing their respective weaknesses. Therefore, it is crucial to explore the question: Can we integrate both symbolic and textual approaches into a hybrid method to enhance tabular reasoning?
@@
In this paper, we introduce H-STAR which integrates symbolic and textual reasoning abilities of LLMs, achieving the best of both worlds. H-STAR decomposes the complex task of table reasoning task into two sub-tasks: a) Table Extraction, where it extracts a question-specific table, and b) Adaptive Reasoning, where it combines SQL-based and text-based reasoning. In the table extraction phase, we employ a step-by-step ‘multi-view’ chain, first identifying relevant columns using both the original and transposed tables. The filtered table is then used for row extraction, contrary to other approaches, which use the original table for both row and column selection in a single view Ye et al. (2023); Nahid and Rafiei (2024).
Focusing on relevant table cells helps LLMs by providing the right context, thereby reducing hallucinations in reasoning. We utilize both SQL-based and text-based techniques in H-STAR’s table extraction phase. In the adaptive reasoning phase, we guide the LLM through a process of step-by-step ‘adaptive’ reasoning, where it chooses between symbolic (SQL) and semantic (textual) methods. H-STAR tailors its reasoning strategies based on the nature of the questions, employing symbolic reasoning for quantitative, mathematical, and logical tasks, while opting for semantic reasoning for direct lookup, common-sense, and challenging lexical queries.
@@
We evaluate the efficacy of H-STAR across three tabular benchmarks involving table QA and table fact verification. Our experiments show that using the combined approach of H-STAR, aligned with multi-view extraction and adaptive reasoning outperforms all prior state-of-the-art approaches. Our results demonstrate that H-STAR is robust and generalizable across LLMs. Through ablations, error analysis, and performance evaluations, we illustrate the effectiveness of H-STAR. Our main contributions are:
@@
(c) Our experiments reveal H-STAR’s efficiency in table extraction using its precise ‘multi-view’ mechanism. Our analysis also highlights H-STAR’s superior reasoning capabilities, emphasizing the necessity of every step.
@@ 2 H-STAR Approach
[…]
@@ 2.1 Table Extraction
Two-dimensional tables consist of columns (attributes) and rows (values). Table extraction involves a two-step reasoning chain: (1) column extraction and (2) row extraction (see Figure 3). As illustrated in Algorithm 1, for a given table T, table extraction returns a table Tcr (Steps 1-8) based on the input query Q.
@@
As shown in Figure 3, H-STAR uses a ‘multi-view’ technique for column extraction, working with both the original table T (Step 1) and its transposed form TT (Step 2). This ensures that the LLM semantically understands and extracts correct attributes regardless of the table structure. Given an input table T and the question Q (‘How long did it take the New York Americans to win the National Cup after 1936?’), the LLM employs colsql to generate a SQL query to extract columns ‘year’ (C1). This is followed by the text-based column verification step coltext on TT returning a new set of columns C2 (‘year’, ‘national cup’). Table T is filtered for columns ‘year’, ‘national cup’ (C) derived from Steps 1 and 2 to obtain a column-filtered table $\textsc{T}_{\textsc{c}}$ (Step 4).
@@
0:  $\hat{A}$rightarrow Predicted AnswerTable ExtractionStep 1: Column Extraction
1:   C1$\leftarrow$ colsql(T, Q) $\triangleright$ SQL extraction
2:   C2 $\leftarrow$ coltext(TT, Q, C1) $\triangleright$ Text $\triangleright$ $\textsc{T}^{\textsc{t}}$ refers to the transposed table
@@
8:   TCR $\leftarrow$ TC.filter(R) $\triangleright$ Filter table $\textsc{T}_{\textsc{c}}$ with rows $\textsc{R}^{\prime}$ to get $\textsc{T}_{\textsc{cr}}$daptive Reasoning
@@
After obtaining the filtered table $\textsc{T}_{\textsc{c}}$ with relevant columns, H-STAR proceeds to the row extraction phase culminating in the query-specific table Tcr (Steps 5-8). The filtered table $\textsc{T}_{\textsc{c}}$ has fewer tokens than T, enabling more efficient row extraction. The SQL query generated by rowsql on the column-filtered table $\textsc{T}_{\textsc{c}}$ extracts the relevant rows R1 (‘row 18’). While it returns the appropriate row, it is not sufficient to answer the question. To overcome these limitations, we use text-based row verification rowtext to obtain the missing rows, ultimately retrieving ‘row 1’ and ‘row 18’ ($\textsc{R}_{2}$). Combining $\textsc{R}_{1}$ and $\textsc{R}_{2}$ gives the relevant rows $\textsc{R}^{\prime}$ (‘row 1’, ‘row 18’), resulting in the final table $\textsc{T}_{\textsc{cr}}$.
@@ 2.2 Adaptive Reasoning
To leverage the strengths of textual reasoning while overcoming its quantitative limitations, we introduce an adaptive reasoning framework. Assessing the nature of the question, the LLM incorporates symbolic reasoning only when required. For instance, in Figure 3, when faced with a question like "How long did it take the New York Americans to win the National Cup after 1936?", LLM requires a count as an answer, and our pipeline prioritizes symbolic reasoning (fsql) through SQL-generated code. This code is executed on a SQL engine, enabling precise answers from the table $\textsc{T}_{\textsc{cr}}$. The output from this SQL-based evaluation Ev serves as additional evidence, enhancing the textual reasoning process (ftext). By explicitly integrating this quantitative approach, our algorithm effectively addresses LLMs’ limitations in handling mathematical operations, thereby improving the accuracy and robustness when tackling questions that demand numerical reasoning.
@@ 3 Experiments
In our research, we use state-of-the-art large language models (LLM) such as PaLM-2 Anil et al. (2023) and GPT-3.5-Turbo Brown et al. (2020); OpenAI (2023) for table reasoning tasks. Our model inputs include in-context examples, the table, and the question for each step of the pipeline. Previous studies, such as Chain-of-Table and TabSQLify, influenced our model choices to ensure a fair comparison. All baselines with PALM-2 are from the propriety PALM-2 model, which is different from the public PALM-2 API (used in H-STAR and TabSQLify).
@@
3.1 Results and Analysis
Table 1 compares the performance of different methods on the table reasoning tasks, TabFact, and WikiTQ datasets, across several LLMs including GPT-3.5-Turbo and PaLM-2. This comparison involves evaluating against generic reasoning, table manipulation techniques, and our method, H-STAR. Appendix A provides results on the FeTaQA dataset in addition to comparisons with other baseline models on the WikiTQ and TabFact dataset.
@@ 3.1 Results and Analysis
H-STAR 83.74 68.85 85.92 68.51
Table 1: Comparison of various methods across datasets. The results are reported from the Chain-of-Table paper for fair comparison. * Chain-of-Table uses a proprietary PaLM-2 model that is better than their publicly available one.
Analysis. (a) On the WikiTQ dataset, H-STAR demonstrates superior performance with both GPT-3.5-Turbo and PaLM-2 models, surpassing all baseline methods. Specifically, H-STAR achieves an accuracy of 68.85% on WikiTQ with GPT-3.5-Turbo, marking a significant improvement of 17.01% over the vanilla GPT-3.5-Turbo model and a 4% increase over the previous best method, TabSQLify. (b) On the TabFact dataset, H-STAR outperforms all methods with GPT-3.5-Turbo, attaining an accuracy of 83.74%, which is a 13.29% improvement over the vanilla model and a 4% margin over the previous best, Chain-of-Table. However, with PaLM-2, H-STAR achieves comparable performance on TabFact while outperforming Chain-of-Table by 1.2% on WikiTQ.
@@
Stage-1 in H-STAR involves the extraction of the table most relevant to the question. Our table extraction method uses a two-step chain to select the relevant columns followed by the rows. Figure 4 showcases the effectiveness of the technique to extract only the relevant information from the table. Figure 4 compares the number of average table cells for the extracted table for H-STAR, H-STAR without row extraction, and H-STAR without column extraction with other baselines.
@@
Figure 4: Comparison of average table cells in the final table.
Performance on Longer Tables.
Table-CoT Chen (2023) highlights that longer table size, with tables exceeding 30 rows, is a significant cause of erroneous generations. This observation is supported by a decline in LLM performance as the number of tokens in the table increases. We classify LLM performance based on total tokens into three groups - small (< 2000 tokens), medium (2000 to 4000 tokens), and large (> 4000 tokens). We then compare H-STAR with other table manipulation methods, including BINDER, DATER, Chain-of-Table, and TabSQLify.
@@
3.2 Efficiency Analysis

### v2

@@ Abstract
Tabular reasoning involves interpreting natural language queries about tabular data, which presents a unique challenge of combining language understanding with structured data analysis. Existing methods employ either textual reasoning, which excels in semantic interpretation but struggles with mathematical operations, or symbolic reasoning, which handles computations well but lacks semantic understanding. This paper introduces a novel algorithm H-STAR that integrates both symbolic and semantic (textual) approaches in a two-stage process to address these limitations. H-STAR employs: (1) step-wise table extraction using ‘multi-view’ column retrieval followed by row extraction, and (2) adaptive reasoning that adapts reasoning strategies based on question types, utilizing semantic reasoning for direct lookup and complex lexical queries while augmenting textual reasoning with symbolic reasoning support for quantitative and logical tasks. Our extensive experiments demonstrate that H-STAR significantly outperforms state-of-the-art methods across three tabular question-answering (QA) and fact-verification datasets, underscoring its effectiveness and efficiency. [^1] The codes and data are available at: https://github.com/nikhilsab/H-STAR
@@ 1 Introduction
Tabular data is one of the most widely used formats for storing structured information in real-world applications. Table-based reasoning presents an inherently challenging problem, requiring logical, mathematical, and textual reasoning over unstructured queries and structured tables Ye et al. (2023). Thus, understanding and inferring tabular data has become a significant area of research in natural language processing. Tabular reasoning tasks (illustrated in Figure 1), such as table-based question answering Pasupat and Liang (2015); Nan et al. (2022) and table-based fact verification Chen et al. (2019), have been extensively explored in the past.
@@
Advancements in Large Language Models (LLMs) have led to better performance across various tasks through carefully crafted prompts. In the domain of table reasoning, symbolic approaches such as Program of Thought Chen et al. (2023) and textual methods like Chain of Thought (CoT) Wei et al. (2022); Brown et al. (2020) have been explored. However, when applied individually (see Figure 2) these methods often struggle due to the complexities imposed by the intricate mix of numerical, temporal, and textual data, coupled with complex table structures Shi et al. (2023); Liu et al. (2023); Sui et al. (2024). Textual reasoning excels in natural language understanding but often misinterprets table structures and struggles with quantitative reasoning. Conversely, SQL-based approaches are strong in quantitative problem-solving but perform poorly on noisy or unstructured inputs Liu et al. (2023). As a result, recent techniques Cheng et al. (2022); Ye et al. (2023); Wang et al. (2023); Nahid and Rafiei (2024) that rely solely on textual or symbolic reasoning suffer from similar limitations. Therefore, it is crucial to explore the question: How can symbolic and textual approaches be integrated into a hybrid method for tabular reasoning that leverages their complementary strengths while mitigating their individual limitations?
@@
In this paper, we introduce H-STAR which combines symbolic and textual reasoning abilities of LLMs, achieving the best of both worlds. H-STAR decomposes the table reasoning task into two sub-tasks: (a) Table Extraction and (b) Adaptive Reasoning. In the table extraction phase, we employ a step-by-step ‘multi-view’ chain, first identifying relevant columns using the original table and its transposed form. The filtered table is then used for row extraction, contrary to other approaches, which use the original table for both row and column selection Ye et al. (2023); Nahid and Rafiei (2024). Focusing on relevant table cells helps LLMs by providing the right context, thereby reducing hallucinations in reasoning. We employ a combination of SQL-based and text-based techniques in H-STAR’s table extraction phase. In the adaptive reasoning phase, we use the language comprehension capabilities of LLMs to guide them with few-shot examples to decide when to support semantic methods with symbolic (SQL) approaches. H-STAR employs semantic reasoning universally, using it exclusively for direct lookup, common-sense, and complex lexical queries while using an additional SQL step for quantitative, mathematical, and logical tasks. This strategy optimizes performance on diverse question types, supporting text comprehension with computation when required.
@@
We demonstrate the efficacy of H-STAR across three tabular benchmarks involving table QA and table fact verification. Our experiments show that using the combined approach of H-STAR, aligned with multi-view extraction and adaptive reasoning outperforms all prior state-of-the-art approaches. Our analysis demonstrates that H-STAR is robust and generalizable across LLMs, showcasing its superior reasoning capabilities across various tasks. Our main contributions are:
@@
(c) Our experiments showcase H-STAR’s efficient table extraction and enhanced reasoning capabilities across different models and datasets, emphasizing the importance of every stage in achieving improved performance.
@@ 2 H-STAR Approach
H-STAR decomposes the table reasoning task into extraction and reasoning stages, combining LLM’s textual comprehension with symbolic reasoning. It converts the original table (T) into a query-specific table (TCR) (refer Algorithm 1). Figure 3 illustrates our H-STAR approach. Unlike, DATER Ye et al. (2023) and TabQSLify Nahid and Rafiei (2024), which use textual reasoning and text-to-SQL respectively, H-STAR integrates both reasoning types in a complementary manner for table extraction. Our reasoning step dynamically uses symbolic techniques alongside semantic methods for quantitative questions, overcoming the limits of pure textual reasoning. H-STAR operates in two main stages: 1) Table Extraction and 2) Adaptive Reasoning.
@@ 2.1 Table Extraction
Two-dimensional tables consist of columns and rows. Table extraction involves a two-step reasoning chain: (1) column extraction and (2) row extraction (see Figure 3). As illustrated in Algorithm 1, for a given table T, table extraction returns a table TCR (lines 1-8) based on the input query Q.
@@
As shown in Algorithm 1, H-STAR uses a ‘multi-view’ technique, first using the original table T (line 1) followed by its transposed form TT (line 2) for column extraction. Changing the ‘view’ acts as a verifier accounting for any information missed due to the table structure, thus improving the overall performance (see Appendix A.4). In Figure 3, given an input table T and the question Q (‘How long did it take the New York Americans to win the National Cup after 1936?’), the LLM employs colsql to generate an SQL query to extract columns ‘year’ (C1). This is followed by the text-based step coltext which processes table TT to return columns C2 (‘year’, ‘national cup’). This step captures the ‘national cup’ information relevant to answer the query, which is missed by colsql. Table T is then filtered for columns ‘year’, ‘national cup’ (C’) derived from C1 and C2 to obtain table Tc (lines 3,4).
@@
0:  $\hat{A}$rightarrow Predicted AnswerStage A: Table ExtractionColumn Extraction
1:   C1$\leftarrow$ colsql(T, Q) $\triangleright$ SQL approach
2:   C2 $\leftarrow$ coltext(TT, Q) $\triangleright$ Text approach $\triangleright$ $\textsc{T}^{\textsc{t}}$ refers to the transposed table
@@
8:   TCR $\leftarrow$ TC.filter(R) $\triangleright$ Filter table $\textsc{T}_{\textsc{c}}$ with rows $\textsc{R}^{\prime}$ to get $\textsc{T}_{\textsc{cr}}$tage B: Adaptive Reasoning
@@
After obtaining the filtered table TC with relevant columns (Algorithm 1, line 4), H-STAR proceeds to the row extraction phase culminating in the query-specific table TCR (lines 5-8). The filtered table TC has fewer tokens than T, fitting within the token limit and enabling more efficient row extraction. As shown in Figure 3, the SQL query generated by rowsql on the column-filtered table TC extracts the relevant rows R1 (‘row 18’). While it returns the appropriate row, it is not sufficient to answer the question. To overcome these limitations, we use text-based row verification rowtext to obtain the missing rows, ultimately retrieving ‘row 1’ and ‘row 18’ (R2). Combining R1 and R2 gives the relevant rows R’ (‘row 1’, ‘row 18’), resulting in the final table TCR.
@@ 2.2 Adaptive Reasoning
We propose an adaptive reasoning framework that harnesses the strengths of textual reasoning while mitigating its quantitative limitations through symbolic reasoning. First, we prompt LLMs to leverage their language understanding capabilities to evaluate the query requirements. We apply symbolic reasoning for queries necessitating quantitative analysis and use the generated output (Ev) along with the table (TCR) as the input to the final textual reasoning step. For instance, in Figure 3, when faced with a question like "How long did it take the New York Americans to win the National Cup after 1936?", LLM needs to calculate the answer, and our pipeline prioritizes symbolic reasoning (fsql) through SQL-generated code. This code is executed on a SQL engine, enabling precise answers from the table TCR. The output from this SQL-based evaluation Ev serves as additional evidence, supporting the final reasoning step (ftext). By explicitly integrating this quantitative approach, our algorithm effectively addresses LLMs’ limitations in handling quantitative reasoning, thereby improving the accuracy and robustness when tackling questions that demand numerical reasoning.
@@ 3 Experiments
In our research, we use state-of-the-art large language models (LLM) such as Gemini-1.5-Flash Reid et al. (2024), PaLM-2 Anil et al. (2023), GPT-3.5-Turbo, GPT-4o-mini OpenAI (2023), and the open-source Llama-3-70B Dubey et al. (2024) for table reasoning tasks. Our model inputs include in-context examples, the table, and the question for each step of the pipeline. All baselines with PALM-2 are from the propriety PALM-2 model, which is different from the public PALM-2 API (used in H-STAR and TabSQLify).
@@
3.1 Main Results
Table 1 compares the performance of different methods on the TabFact, and WikiTQ datasets, across GPT-3.5-Turbo and PaLM-2. This comparison involves evaluating against generic reasoning, table manipulation techniques, and H-STAR. Appendix A includes analyses on the FeTaQA dataset.
@@ 3.1 Main Results
H-STAR 85.03 69.56 86.51 68.62
Table 1: Comparison of various methods across datasets. The results are reported from the Chain-of-Table paper for fair comparison. * Chain-of-Table uses a proprietary PaLM-2 model that is better than the publicly available version.
Analysis. (a) On WikiTQ dataset, H-STAR surpasses all baselines with both GPT-3.5-Turbo and PaLM-2 models. H-STAR achieves an accuracy of 69.56% on WikiTQ with GPT-3.5-Turbo, marking an improvement of 17.72% over the vanilla GPT-3.5-Turbo model. (b) On the TabFact dataset, H-STAR outperforms all methods with GPT-3.5-Turbo, attaining an accuracy of 85.03%, which is a 14.58% improvement over the vanilla model. With PaLM-2, H-STAR achieves comparable performance with Chain-of-Table on TabFact while improving over it by 1.3% on WikiTQ.
Comparison Across Methods
Table 2 compares H-STAR with ReAcTable Zhang et al. (2023) and SYNTQA Zhang et al. (2024b), which use both textual and symbolic approaches for table reasoning. SYNTQA (GPT) uses the GPT model to decide between semantic and symbolic approaches, applying the chosen method for reasoning based on its selection. Additionally, it is compared with ALTER Zhang et al. (2024a), which augments both queries and table data to facilitate reasoning.
TabFact WikiTQ
ReAcTable 73.1 52.5
ALTER 84.3 67.4
SYNTQA (GPT) - 65.2
H-STAR 85.0 69.6
Table 2: Comparison of various methods across datasets on GPT-3.5-Turbo.
Analysis. H-STAR outperforms methods that use symbolic and textual approaches, such as ReAcTable, SYNTQA(GPT), and ALTER, showing the effectiveness of its complementary reasoning in a two-stage process. Appendix A.1 presents a comparison with more methods.
Performance Across LLMs
Table 3 compares the performance of more advanced models like Gemini-1.5-Flash, GPT-4o-mini, and the open-source Llama-3-70B on TabFact and WikiTQ datasets. The results emphasize H-STAR’s generalizability across different models and consistent improvements across datasets. Even advanced models show significant benefits by using H-STAR.
GPT-4o-mini Gemini-1.5 Llama-3
TF WTQ TF WTQ TF WTQ
Generic Reasoning
End-to-End QA 73.22 59.43 81.12 58.47 78.41 57.89
CoT 75.99 64.31 79.99 64.11 75.34 65.49
Table Manipulation
TabSQLify 78.30 68.74 79.50 63.92 60.70 66.85
Chain-of-Table 85.09 68.53 86.95 70.05 85.86 70.76
H-STAR 89.42 74.93 89.08 73.14 89.23 75.76
Table 3: Comparison of various models across datasets. TF: TabFact; WTQ: WikiTQ
Analysis. (a) H-STAR shows much higher accuracy for GPT-4o-mini, scoring 89.42% on TabFact and 74.93% on WikiTQ, compared to the End-to-End QA and Chain of Thought methods, which achieve 73.22% and 75.99% on TabFact and 59.43% and 64.31% on WikiTQ. This marks a 16.2% improvement on TabFact and 15.5% on WikiTQ. (b) For Gemini-1.5-Flash, accuracy rises from 81.12% and 58.47% respectively to 89.08% on TabFact and 73.14% on WikiTQ. (c) Similarly, for Llama-3-70B, H-STAR improves the performance across both the datasets by a margin of 11% on TabFact and 18% on WikiTQ, showing consistent gains across models and datasets.
3.2 Efficiency Analysis
@@
Stage-1 in H-STAR involves the extraction of the table most relevant to the question. Our table extraction method uses a two-step chain to select the relevant columns followed by the rows. Figure 4 compares the number of average table cells for the extracted table for H-STAR, H-STAR without row extraction, and H-STAR without column extraction with other baselines.
Figure 4: Comparison of average table cells in the final table.
@@
Efficiency on Longer Tables.
Table-CoT Chen (2023) highlights that longer table size, with tables exceeding 30 rows, is a significant cause of erroneous generations. This observation is supported by a decline in LLM performance as the number of tokens in the table increases. In Table 4, we classify LLM performance based on total tokens into three groups - small (< 2000 tokens), medium (2000 to 4000 tokens), and large (> 4000 tokens) and compare H-STAR with table manipulation like methods BINDER, DATER, Chain-of-Table, and TabSQLify.
@@
3.3 Error Analysis
H-STAR
The disjoint, step-wise nature of H-STAR enables identifying and analyzing failures. Figure 5 shows 100 randomly selected instances from the Tabfact and WikiTQ datasets where H-STAR returns incorrect answers. In this study, ‘Missing Columns’ and ‘Missing Rows’ refer to missing necessary columns and rows, respectively. ‘Incorrect Reasoning’ occurs when H-STAR extracts the correct table, but LLM fails to produce the correct answer. ‘Incorrect Annotations’ include semantically identical answers in different formats, ambiguous questions, and incorrect gold answers. Figure 5 shows that for 100 TabFact failures, 2% are missing columns, 9% missing rows, 79% incorrect LLM reasoning, and 10% incorrect annotations. For WikiTQ, 6% are missing columns, 17% missing rows, 54% incorrect reasoning, and 23% incorrect annotations.
Figure 5: Error distribution on 100 error samples across datasets for H-STAR (GPT-3.5-Turbo).
Analysis. Fewer errors in column and row extraction indicate that H-STAR effectively retrieves the necessary table data. This improvement allows us to "shift left" by addressing issues earlier in the pipeline, enhancing overall performance. The higher percentage of errors in reasoning does not reflect poorly on the LLM; instead, it underscores the effectiveness of our table extraction process.
H-STAR vs Others
Figure 6 compares errors in H-STAR, with TabSQLify, and BINDER on 100 WikiTQ samples where TabSQLify fails. TabSQLify extracts wrong tables for 62 out of 100 (6 missing columns, 56 missing rows). Out of the 38 samples that remain, it incorrectly reasons on 29 of the samples (76%). BINDER does not handle table extraction and instead relies on generating multiple neural SQL queries for final reasoning. However, it fails to reason correctly in 70 out of 100 instances.
Figure 6: Analysis of error types in 100 samples from WikiTQ where TabSQLify fails. Note: As we move from left to right, the total samples decrease for each stage in the pipeline.
In contrast, H-STAR proves more effective in both table extraction and reasoning, generating correct answers on 43 samples. Additionally, it demonstrates fewer errors in table extraction (4 missed columns; 16 missed rows) and reasoning (27 of the remaining 80 samples, i.e. 34%).
4 Ablation Study
To evaluate the significance of the two primary stages in H-STAR: a) Table extraction and b) Adaptive reasoning, we conducted an ablation study, removing one stage at a time. For the first condition, the final adaptive reasoning stage was substituted with a basic chain-of-thought reasoning process, while maintaining the table extraction stage. Conversely, the second condition retained the adaptive reasoning stage but omitted the table extraction stage, specifically by removing the row extraction and column extraction steps individually and then collectively. A detailed assessment of the contribution of each stage to the overall performance of the method is shown in Table 5. It highlights the importance of each step in the pipeline. Removing any step can result in a performance loss especially adaptive reasoning which causes a 7% drop for both TabFact and WikiTQ datasets. For a detailed error case study and further analysis, refer to the Appendix E.
Method TabFact WikiTQ
H-STAR 86.51 68.62
w/o row extraction 86.17 66.30
w/o column extraction 84.04 67.03
w/o table extraction 83.79 63.58
w/o adaptive reasoning 79.35 61.47
Table 5: Performance on TabFact and WikiTQ datasets.
4.1 Impact of the Hybrid Approach
We perform a quantitative study of the impact of the hybrid approach in H-STAR. Figures 16, 17 (Appendix E) depict how using only one of the approaches results in insufficient outputs for column and row extraction tasks respectively. Furthermore, Table 6 shows that not employing a hybrid approach results in a significant performance loss.
Method TabFact WikiTQ
CoT 79.05 60.43
H-STAR ${}_{\text{text}}$ 79.43 61.47
Text-to-SQL 52.05 42.03
H-STAR ${}_{\text{sql}}$ 58.50 46.09
H-STAR 86.51 68.62
Analysis. SQL methods though efficient with numerical reasoning, are highly sensitive to data variations like irregular formatting and mixed data types, leading to a performance drop. In contrast, textual reasoning is more resilient, offering fuzzy matching and better interpretation of data. By combining textual and symbolic reasoning, H-STAR uses their complementary strengths, achieving superior performance.
4.2 Symbolic vs Semantic Approaches
As shown in Table 7 removing SQL extraction drops the accuracy to 85.22% on TabFact and 64.39% on WikiTQ, removing text extraction causes a drop to 83.74% and 60.31%, respectively. SQL reasoning removal results in 84.48% and 64.76%, but omitting text reasoning causes the largest drop, to 58.70% and 54.35%.
Method TabFact WikiTQ
H-STAR 86.51 68.62
w/o SQL extraction 85.22 64.39
w/o text extraction 83.74 60.31
w/o SQL reasoning 84.48 64.76
w/o text reasoning 58.70 54.35
Table 7: Performance of H-STAR after systematically removing symbolic part and semantic parts from: 1) table extraction (both row and column extraction); 2) adaptive reasoning.
Analysis. Table 7 shows that text-based approaches excel in both table extraction and reasoning. The evaluation datasets often contain noisy data that SQL-based approaches struggle with due to their reliance on structured schemas. Since text-based reasoning methods are more effective at handling such irregularities, they achieve reasonably high scores, albeit lower than our hybrid approach.
5 Key Findings
Firstly, our evaluation demonstrates that our hybrid approach achieves substantial improvements, surpassing the performance of previous state-of-the-art methods across various table reasoning tasks. Secondly, the quantitative analysis demonstrates that our ‘multi-view’ approach extracts tables specific to the query. Furthermore, the qualitative analysis highlights fewer errors in our table extraction method compared to prior approaches, confirming a decrease in irrelevant information.
Thirdly, our analysis indicates consistent performance even with longer tables, emphasizing our method’s effectiveness in accurately extracting relevant information by filtering out noise. Lastly, the ablation study shows that decomposing the task into sub-tasks significantly enhances the overall performance, with each sub-task playing a crucial role in achieving superior results. Moreover, it also highlights the constraints of relying solely on either text or SQL approaches, which are effectively addressed by our H-STAR approach. Together, these findings emphasize the substantial advantages of our hybrid and multi-view approach in addressing complex table reasoning tasks.
6 Related Work
Decomposing problems into smaller, manageable tasks has proven effective in solving complex reasoning challenges Zhou et al. (2022); Khot et al. (2022). Recent techniques in table reasoning follow this approach, either by breaking tasks into fixed sub-tasks Cheng et al. (2022); Ye et al. (2023); Nahid and Rafiei (2024) or by employing iterative methods Jiang et al. (2023); Zhang et al. (2023); Wang et al. (2023). BINDER Cheng et al. (2022) is an SQL-based approach that modifies SQL statements to include LLM API calls within SQL statements. ALTER Zhang et al. (2024a) augments both the queries along with the table data. DATER Ye et al. (2023) and TabSQLify Nahid and Rafiei (2024) are table decomposition methods that use semantic and symbolic reasoning respectively. Chain-of-Table Wang et al. (2023) uses a textual reasoning approach to update tables iteratively before the final reasoning step. ReAcTable Zhang et al. (2023) extends the ReAct Yao et al. (2023) framework to table reasoning by employing step-by-step reasoning, iteratively generating sub-tables using external tools like SQL and Python. A detailed explanation has been provided in Appendix C. SYNTQA Zhang et al. (2024b) is an ensemble approach employing an answer selection mechanism selecting between text-to-SQL and text-based models. In contrast, H-STAR follows a fixed two-stage table extraction and reasoning pipeline, applying both symbolic and semantic approaches complementarily at each stage, which leads to improved performance.
7 Conclusion
Limitations
Ethics Statement
Acknowledgments
Appendix A Supplementary Analysis
A.1 Comparison With Diverse Methods
Tables 8 and 9 show results of other baseline methods on WikiTQ and TabFact datasets respectively. Our method, H-STAR, outperformed all other models by substantial margins.
Model Method Accuracy
Pre-trained/fine-tuned TaPas 48.8
GraPPa 52.7
LEVER 62.9
ITR 63.4
SYNTQA(GPT)* 70.4
TableCoT 48.8
DATER 65.9
BINDER 61.9
ReAcTable 65.8
GPT-3.5-Turbo ReAcTable 52.5
TableCoT 52.4
StructGPT 52.2
ALTER 67.4
SYNTQA(GPT)* 65.2
GPT-3.5-Turbo H-STAR 69.6
Table 8: Comparison of H-STAR with additional methods on WikiTQ dataset. * SYNTQA (GPT) uses GPT-3.5-Turbo to choose between SQL/text reasoning and uses: 1) fine-tuned models for reasoning; 2) GPT-3.5-Turbo for reasoning.
Model Method Accuracy
Pre-trained/fine-tuned Table-BERT 68.1
LogicFactChecker 74.3
SAT 75.5
TaPas 83.9
TAPEX 85.9
SaMoE 86.7
PASTA 90.8
Codex TableCoT 72.6
DATER 85.6
BINDER 85.1
ReAcTable 83.1
GPT-3.5-Turbo ReAcTable 73.1
TableCoT 73.1
ALTER 84.3
GPT-3.5-Turbo H-STAR 85.0
A.2 FeTaQA
Prompting ROUGE-1 ROUGE-2 ROUGE-L
PaLM-2*
DATER 0.63 0.41 0.53
Chain-of-Table 0.66 0.44 0.56
GPT-3.5-Turbo
Table-CoT 0.62 0.39 0.51
TabSQLify 0.58 0.35 0.48
H-STAR 0.62 0.39 0.52
Table 10: Performance on FeTaQA dataset. PaLM-2* refers to the Google proprietary PaLM-2 model.
Human Evaluation
Method Fluent Correct Adequate Faithful
T5-large 94.6 54.8 50.4 50.4
Human 95 92.4 95.6 95.6
TableCoT 96 82 75 87
TabSQLify 97 88 84 93
H-STAR 96.6 87.6 89.6 94
Table 11: Human evaluation results on FeTaQA dataset.
Qualitative Analysis
In this section, we perform a qualitative analysis on 100 samples of FeTaQA with TabSQLify Nahid and Rafiei (2024). The study aims to compare the generations of H-STAR with TabSQLify by evaluating the answer quality. The answers are evaluated based on their fluency and adequacy when compared with the gold answers. Table 12 shows that H-STAR has better quality generations in 26% samples, as opposed to only 18% for TabSQLify. ‘Both’ indicates samples where both the algorithms perform on par. This can be attributed to the use of the same LLM. Many failures for TabSQLify are derived from an insufficient table, which further highlights the limitations of using a single-view approach for table extraction.
Method % of samples
Both 56%
TabSQLify 18%
H-STAR 26%
Table 12: Comparison of H-STAR with TabSQLify on 100 samples.
Figure 8 showcases such an example where TabSQLify returns ‘information not provided’ because of incorrect of an extracted table. It can be noted that H-STAR generates an answer comparable with the ground truth for the same. Figure 9 visually depicts the outputs from both methods. It illustrates the evaluation process undertaken to compare the quality of the outputs.
Figure 9: Comparison of the outputs with TabSQLify.
A.3 Analysis: Generated Samples